2NOBr(g) 2NO(g) + Br2(g) 


If 0.364 moles of NOBr(g)0.448 moles of NO, and 0.543 moles of Br2 are at equilibrium in a 19.1 L container at 464 K, the value of the equilibrium constant, Kc, is  .


1
Expert's answer
2022-05-09T04:09:04-0400

Solution:

Balanced chemical equation:

2NOBr(g) → 2NO(g) + Br2(g) 


The expression for the equilibrium constant Kc:



[Br2] = (0.543 mol) / (19.1 L) = 0.02843 M

[NO] = (0.448 mol) / (19.1 L) = 0.02346 M

[NOBr] = (0.364 mol) / (19.1 L) = 0.01906 M

Thus,

Kc = (0.02843) × (0.02346)2 / (0.01906)2 = 0.043

Kc = 0.043 M


Answer: The value of the equilibrium constant (Kc) is 0.043